OpenStudy (anonymous):

let p be some fixed real number. find the linearization of f(x) = (1 + x)^p at 0

OpenStudy (anonymous):

y = (1+x)^p ln(y) = p*ln(1+x) Say now Y = ln(y) and X = ln(1+x), then: Y = p*X That is a linear function now

OpenStudy (anonymous):

is that kind of linearization you want?

OpenStudy (anonymous):

ohh i guess you want the linearization by taylor series, right?

OpenStudy (anonymous):

\[f(x) \approx \sum_{n=0}^{\infty} \frac{f^n(0) \cdot (x)^n}{n!}\]

OpenStudy (anonymous):

that is the Maclaurin series, which is an particular case of taylor series, just apply the formula till the first order

OpenStudy (anonymous):

Okay, let's make the linearization of f(x) at x=a: f(x) = f(a) + f'(a)(x-a), for any "a" in the range of f(x) f'(a) is the derivative of f(x) avaluated at x=a

OpenStudy (anonymous):

yes, the tangent line as in linear approximation

OpenStudy (anonymous):

I have L(x) = f(0) + p(1+x)^(p-1) (0-p)

OpenStudy (anonymous):

good catch

OpenStudy (anonymous):

it should be: L(x) = f(0) + p(1+0)^(p-1) (x-0)

OpenStudy (anonymous):

simplifying: L(x) = f(0) + px